Three consecutive numbers have a sum of 33. What is the middle number?
Found 2 solutions by Alan3354, jim_thompson5910:
Answer by Alan3354(69443)

33/3 is the middle number

Answer by jim_thompson5910(35256)

x + (x+1) + (x+2) = 33

x + x+1 + x+2 = 33

3x + 3 = 33

3x = 33-3

3x = 30

x = 30/3

x = 10

The first number is 10.

x+1 = 10+1 = 11.

The second number is 11.

x+2 = 10+2 = 12

The third number is 12.

So the middle number is 11.